OpenStudy (anonymous):

a recipe that makes 9 servings calls for 4/5 cup of orange and 2/3 cup of lemonade. jeff modifies the recipe so that it can serve 10. how many cups of orange does jeff need? a. 7 1/5 b. 8/9 c. 20/27 d. 4/45

OpenStudy (anonymous):

by ratio and proportion\[\frac{9~servings}{\frac{4}{5}~orange}=\frac{10~servings}{x~orange}\]what do you think? @lbjm